# A flexible class of latent variable models for the analysis of antibody response data

### v1

@@ 5.1 Modelling apical membrane antigen 1 (AMA1) concentrations
The results of this step are presented in Figure 4. As age increases, the distribution of AMA1 antibody concentrations shifts toward higher values, indicating an overall increase in malaria exposure and infection. Another noteworthy pattern is that the bimodality, or apparent separation between the two mixture components, becomes less distinct above 15 years of age. Examining the estimates of the Beta shape parameters, $\alpha$ and $\beta$, we also observe an overall increase with age, with most of the estimate values for both parameters below 1 (indicating a U-shaped Beta) for ages below 15 years. This supports the initial interpretation based on the histograms: at younger ages, the population is more polarized between healthy and unhealthy states, whereas with increasing age the fitted Beta distributions transition gradually from a U-shaped to a unimodal form. The increasing values of $\alpha$ and $\beta$ across age groups lead to a mean of $T$ that peaks in the $[15,20)$ interval and then gradually declines, suggesting a mild waning of antibody concentrations at older ages.
Based on the data, individuals below 15 years of age display a distinct immunological profile and antibody dynamics compared with older age groups. To illustrate how this can be accommodated, we adopt a joint modelling framework that accounts for these differences. For individuals under 15, we use a mechanistic representation of $T$ based on the reversible catalytic model (11), capturing the biological processes of antibody acquisition and boosting. For those aged 15 and above, we adopt a more empirical specification that flexibly describes the observed antibody distributions without imposing strong mechanistic assumptions. This represents one of several possible analyses within our framework, not necessarily yielding the best-fitting model, but serves to demonstrate how joint modelling across ages can make fuller use of the available data, avoiding the common practice of dichotomizing or fitting separate models by age group.
@@
This constraint prevents an artificial discontinuity in the age-dependent mean of $T$ and ensures a biologically coherent transition between the early-life dynamics of antibody acquisition and the slower antibody fluctuations observed at later ages when the model is fitted jointly. The change point $\tau$ is estimated simultaneously with the other model parameters, allowing the data to determine the optimal age at which the mechanistic representation transitions to the more flexible empirical specification.
Based on the constraints imposed on the parameters of the Beta mixture model for $T$ (see Section 3.2), an initial fit indicated that $\alpha_{1}$, $\beta_{1}$, $\beta_{2}$, and $p_{0}$ had estimates converging towards their boundary values. The fitted values were close to 1 for $\alpha_{1}$, $\beta_{2}$, and $p_{0}$, while $\beta_{1}$ resulted in a very large value of approximately $2990$. These results suggest that the first mixture component can be approximated with a Dirac measure at zero, representing individuals with very low or no seroreactivity. In other words, this finding motivates a simplified formulation in which the first component of $T$ is treated as a degenerate mass at 0, corresponding to individuals with absent or waning antibody levels, while the second component remains a continuous Beta distribution with $\alpha_{2}$ to be estimated and $\beta_{2}=1$, representing individuals with detectable antibody responses arising from exposure to AMA1. The resulting specification for individuals below $\tau$ years of age thus becomes
$T\sim(1-\pi(a))\,\delta_{0}+\pi(a)\,\mathrm{Beta}(\alpha_{2},1),\qquad\pi(a)=1-\exp(-\lambda a).$ (13)
@@
Table 1 summarises the parameter estimates and uncertainty from the parametric bootstrap procedure for the joint model fitted across the full age range. The estimated change point for the transition from the mechanistic formulation to the regression-based formulation is $\hat{\tau}=20.8$ years. This suggests that the mechanistic framework remains appropriate up to approximately age 20, allowing for a substantially broader age range to be used in transmission intensity estimation than is typically employed in practice. This finding demonstrates that rather than imposing an a priori age restriction, as is sometimes done in serology studies (e.g. yman2016), our model provides an empirical, data-driven approach to determining the appropriate age range for mechanistic modelling.
@@
The left panel of Figure 5 shows the results of the validation procedure described in Section 4.1. The main discrepancy between the fitted model and the observed data occurs in the age group $[6,10)$, where the model produces a distribution with a slightly higher concentration of individuals with elevated antibody levels than observed. For the remaining age groups, the observed deviations are minor. Such differences are expected because the model imposes a relatively strong structure on the age-dependent dynamics to ensure the interpretability of its components. As is often the case, a more empirical specification could yield a closer fit to the data, but at the cost of losing biological interpretability.
@@ 5.2 Modelling merozoite surface protein 1 (MSP1) concentrations
Unlike the analysis for AMA1, here we adopt a more data-driven strategy that fully exploits the flexibility of the latent-variable framework. As we will show, this enables a closer empirical fit to the MSP1 data than in the model used for AMA1, where we instead adopted a more mechanistically motivated approach. Based on the insights from Figure 6, we formulate an age-dependent model in which $T$ follows a Beta distribution with a smoothly varying $\alpha(a)$ and a piecewise age effect for $\beta(a)$:
@@
Table 2 summarises the parameter estimates for the MSP1 model, which point to a more durable and persistent antibody response than that observed for AMA1. The estimated exponents $\gamma_{1}$ and $\delta_{1}$ show that both $\alpha(a)$ and $\beta(a)$ increase with age, but that $\alpha(a)$ grows more rapidly, producing a steady rise in the mean latent immune state $\mathbb{E}[T]$ as given by equation (14). This rise becomes even steeper after the change point at $\zeta=11.6$ years, as reflected by the negative value of $\delta_{2}$, which slows the growth of $\beta(a)$ and therefore increases $\mathbb{E}[T]$ more sharply. The validation results in Figure 7 indicate that the model provides a good overall fit, with only minor deviations of the model-based histograms from the empirical histograms observed across the different age groups.
To interpret these parameter estimates in terms of the observed antibody response, Figure 8 displays the expected mean of $Y$ as a function of age for both MSP1 and AMA1. As anticipated, AMA1 rises sharply in early childhood but then plateaus and declines during adolescence, reflecting the higher individual-level variability in AMA1 antibody dynamics, with rapid boosting and waning following each infection (akpogheneta2008; wipasa2010; ondigo2014). However, the peak around age 20 years may also reflect the higher prevalence of asymptomatic infections with detectable parasitemia in this age range. At this stage, individuals have typically acquired clinical immunity (protection against disease symptoms) but have not yet fully developed anti-parasite immunity, sustaining infections with parasite densities high enough to continue boosting antibody levels. In contrast, MSP1 shows a steady and persistent increase across the full age range (lower panel of Figure 8), consistent with the more consistent dynamics of MSP1 responses and their accumulation through repeated exposure (wu2020; ondigo2014).
We also observe that the MSP1 trajectory exhibits a step change increase at approximately 11 years, captured by the change-point parameter $\zeta$. While the precise mechanism underlying this shift is unclear, plausible explanations include cumulative exposure effects associated with school attendance, immune maturation during early adolescence, or cohort effects reflecting temporal changes in malaria transmission intensity. Distinguishing among these possibilities would require longitudinal data or additional historical information on transmission patterns and intervention coverage.

### v2

@@ 4.2 Model validation
5 Simulation study
To evaluate the performance of the $L_{2}$ histogram-based estimator relative to MLE, we conducted a comprehensive simulation study under different scenarios that mimic specific epidemiological settings. The simulation study aims to compare the two methods in terms of both statistical efficiency (bias and root mean squared error) and computational efficiency (computation time).
For each simulation replicate, we generated data from the latent Beta model specified in (2). Specifically, for each individual $i=1,\ldots,n$, we sampled the latent immune state from a Beta distribution, $T_{i}\sim\text{Beta}(\alpha,\beta)$, and then, conditional on $T_{i}=t_{i}$, sampled the observed antibody concentration from the Gaussian distribution of $Y_{i}\mid T_{i}=t_{i}$.
We considered four scenarios representing distinct epidemiological settings, chosen to reflect the range of serological patterns commonly observed in practice. Each scenario is defined by a specific combination of parameters $(\mu_{0},\mu_{1},\sigma_{0},\sigma_{1},\alpha,\beta)$, with the Beta shape parameters $(\alpha,\beta)$ determining the distribution of latent immune states in the population.
  • The first scenario, bimodal (BM) ($\alpha=0.5$, $\beta=0.5$), uses a strongly U-shaped Beta distribution with equal probability mass near both extremes, corresponding to highly polarized populations with clear separation between seronegative and seropositive groups, with $\mathbb{E}[T]=0.5$.
  • The second scenario, high transmission (HT) ($\alpha=3.0$, $\beta=0.5$), uses a U-shaped Beta distribution concentrated near $T=1$, characteristic of endemic transmission where repeated exposure is common; most individuals show high seroreactivity, with $\mathbb{E}[T]=0.86$.
  • The third scenario, intermediate transmission (IT) ($\alpha=2.0$, $\beta=2.0$), employs a symmetric unimodal Beta distribution with $\mathbb{E}[T]=0.5$, representing populations with balanced exposure patterns and capturing settings with moderate, stable transmission or heterogeneous exposure across subgroups.
  • The fourth scenario, low transmission (LT) ($\alpha=0.5$, $\beta=3.0$), uses a U-shaped Beta distribution concentrated near $T=0$, representing populations with minimal ongoing pathogen exposure, such as post-elimination settings or areas with effective vector control. Under this configuration, the majority of individuals exhibit low seroreactivity, with $\mathbb{E}[T]=0.14$.
For all scenarios, we fixed the conditional distribution parameters at $\mu_{0}=-3.0$, $\mu_{1}=1.0$, $\sigma_{0}=0.8$, and $\sigma_{1}=0.3$, except for the bimodal scenario where we used $\mu_{0}=-3.5$, $\mu_{1}=1.5$, $\sigma_{0}=0.7$, and $\sigma_{1}=0.2$ to increase the separation between low and high seroreactivity states. These parameter values were chosen to reflect realistic antibody concentration ranges on the log scale, with $\mu_{1}-\mu_{0}=4$ (or 4.5 for the bimodal scenario) providing sufficient separation to distinguish the two extremes while $\sigma_{1}<\sigma_{0}$ captures the reduced variability typically observed near assay saturation.
We evaluated performance across six sample sizes: $n\in\{100,500,1000,5000\}$. This range enables us to assess how the relative performance of $L_{2}$ and MLE changes with increasing information. For each combination of scenario and sample size, we generated 1,000 independent replicates. For each simulated dataset, we applied both the $L_{2}$ histogram-based estimator and MLE, recording parameter estimates and computation times for both methods. We construct histograms using Sturges’ rule to determine the number of bins: $k=\lceil\log_{2}(n)+1\rceil$. This binning rule yields between 8 bins ($n=100$) and 14 bins ($n=5000$).
For the estimates of each of the model parameters performance was evaluated using the median bias and the median absolute error. To assess the computational efficiency, we recorded the elapsed wall-clock time (in seconds) required to obtain parameter estimates for each method on every simulation replicate.
5.1 Results
Table 1: Median bias of parameter estimates across scenarios and sample sizes.
Scenario Parameter $n=100$ $n=500$ $n=1000$ $n=5000$
$L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE
BM $\alpha$ 0.105-0.105 0.054-0.054 0.026-0.026 0.008-0.008 0.023-0.023 0.001-0.001 0.005-0.005 0.000-0.000
$\beta$ 0.073-0.073 0.038-0.038 0.028-0.028 0.006-0.006 0.025-0.025 0.001-0.001 0.007-0.007 0.002-0.002
$\mu_{0}$ $0.112$ $0.064$ $0.055$ $0.014$ $0.038$ 0.002-0.002 $0.006$ 0.001-0.001
$\mu_{1}$ 0.118-0.118 0.022-0.022 0.028-0.028 0.002-0.002 0.017-0.017 0.001-0.001 0.001-0.001 0.000-0.000
$\sigma_{0}$ $0.032$ 0.044-0.044 $0.036$ 0.004-0.004 $0.031$ 0.009-0.009 $0.015$ 0.001-0.001
$\sigma_{1}$ $0.060$ 0.023-0.023 $0.084$ 0.004-0.004 $0.076$ 0.003-0.003 $0.061$ 0.000-0.000
HT $\alpha$ 2.105-2.105 2.026-2.026 1.320-1.320 0.494-0.494 0.742-0.742 0.113-0.113 0.027-0.027 $0.269$
$\beta$ 0.286-0.286 0.214-0.214 0.109-0.109 0.071-0.071 0.020-0.020 0.036-0.036 $0.004$ $0.012$
$\mu_{0}$ $1.539$ $0.995$ $1.062$ $0.109$ $0.806$ 0.198-0.198 $0.120$ 0.295-0.295
$\mu_{1}$ 0.052-0.052 0.055-0.055 0.053-0.053 0.019-0.019 0.036-0.036 0.005-0.005 0.009-0.009 $0.011$
$\sigma_{0}$ 0.360-0.360 0.579-0.579 $0.090$ 0.115-0.115 $0.142$ 0.164-0.164 $0.086$ 0.061-0.061
$\sigma_{1}$ $0.013$ 0.015-0.015 $0.022$ 0.005-0.005 $0.016$ 0.002-0.002 $0.019$ 0.003-0.003
IT $\alpha$ 1.826-1.826 1.751-1.751 1.766-1.766 1.534-1.534 1.660-1.660 1.174-1.174 0.850-0.850 0.239-0.239
$\beta$ 1.705-1.705 1.523-1.523 1.612-1.612 1.124-1.124 1.322-1.322 0.696-0.696 0.469-0.469 0.071-0.071
$\mu_{0}$ $1.335$ $1.325$ $1.174$ $1.043$ $1.101$ $0.873$ $0.685$ $0.167$
$\mu_{1}$ 0.747-0.747 0.651-0.651 0.698-0.698 0.396-0.396 0.581-0.581 0.283-0.283 0.234-0.234 0.029-0.029
$\sigma_{0}$ $0.037$ $0.028$ $0.132$ $0.081$ $0.124$ $0.072$ $0.094$ $0.026$
$\sigma_{1}$ 0.036-0.036 0.108-0.108 0.018-0.018 0.131-0.131 $0.062$ 0.059-0.059 $0.034$ 0.022-0.022
LT $\alpha$ 0.444-0.444 0.443-0.443 0.442-0.442 0.411-0.411 0.422-0.422 0.328-0.328 0.306-0.306 0.224-0.224
$\beta$ 2.797-2.797 2.605-2.605 2.828-2.828 2.531-2.531 2.730-2.730 2.347-2.347 2.274-2.274 1.794-1.794
$\mu_{0}$ $0.232$ $0.315$ $0.283$ $0.291$ $0.269$ $0.218$ $0.194$ $0.143$
$\mu_{1}$ 1.777-1.777 1.569-1.569 1.681-1.681 1.426-1.426 1.481-1.481 1.304-1.304 1.403-1.403 1.218-1.218
$\sigma_{0}$ 0.026-0.026 $0.008$ $0.025$ $0.027$ $0.030$ $0.022$ $0.030$ $0.015$
$\sigma_{1}$ 0.030-0.030 0.110-0.110 $0.204$ 0.009-0.009 $0.259$ $0.137$ $0.364$ $0.259$
Table 2: Median Absolute Error (MAE) of parameter estimates across scenarios and sample sizes.
Scenario Parameter $n=100$ $n=500$ $n=1000$ $n=5000$
$L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE
BM $\alpha$ $0.105$ $0.054$ $0.026$ $0.008$ $0.023$ $0.001$ $0.005$ $0.000$
$\beta$ $0.073$ $0.038$ $0.028$ $0.006$ $0.025$ $0.001$ $0.007$ $0.002$
$\mu_{0}$ $0.112$ $0.064$ $0.055$ $0.014$ $0.038$ $0.002$ $0.006$ $0.001$
$\mu_{1}$ $0.118$ $0.022$ $0.028$ $0.002$ $0.017$ $0.001$ $0.001$ $0.000$
$\sigma_{0}$ $0.032$ $0.044$ $0.036$ $0.004$ $0.031$ $0.009$ $0.015$ $0.001$
$\sigma_{1}$ $0.060$ $0.023$ $0.084$ $0.004$ $0.076$ $0.003$ $0.061$ $0.000$
HT $\alpha$ $2.105$ $2.026$ $1.320$ $0.494$ $0.742$ $0.113$ $0.027$ $0.269$
$\beta$ $0.286$ $0.214$ $0.109$ $0.071$ $0.020$ $0.036$ $0.004$ $0.012$
$\mu_{0}$ $1.539$ $0.995$ $1.062$ $0.109$ $0.806$ $0.198$ $0.120$ $0.295$
$\mu_{1}$ $0.052$ $0.055$ $0.053$ $0.019$ $0.036$ $0.005$ $0.009$ $0.011$
$\sigma_{0}$ $0.360$ $0.579$ $0.090$ $0.115$ $0.142$ $0.164$ $0.086$ $0.061$
$\sigma_{1}$ $0.013$ $0.015$ $0.022$ $0.005$ $0.016$ $0.002$ $0.019$ $0.003$
IT $\alpha$ $1.826$ $1.751$ $1.766$ $1.534$ $1.660$ $1.174$ $0.850$ $0.239$
$\beta$ $1.705$ $1.523$ $1.612$ $1.124$ $1.322$ $0.696$ $0.469$ $0.071$
$\mu_{0}$ $1.335$ $1.325$ $1.174$ $1.043$ $1.101$ $0.873$ $0.685$ $0.167$
$\mu_{1}$ $0.747$ $0.651$ $0.698$ $0.396$ $0.581$ $0.283$ $0.234$ $0.029$
$\sigma_{0}$ $0.037$ $0.028$ $0.132$ $0.081$ $0.124$ $0.072$ $0.094$ $0.026$
$\sigma_{1}$ $0.036$ $0.108$ $0.018$ $0.131$ $0.062$ $0.059$ $0.034$ $0.022$
LT $\alpha$ $0.444$ $0.443$ $0.442$ $0.411$ $0.422$ $0.328$ $0.306$ $0.224$
$\beta$ $2.797$ $2.605$ $2.828$ $2.531$ $2.730$ $2.347$ $2.274$ $1.794$
$\mu_{0}$ $0.232$ $0.315$ $0.283$ $0.291$ $0.269$ $0.218$ $0.194$ $0.143$
$\mu_{1}$ $1.777$ $1.569$ $1.681$ $1.426$ $1.481$ $1.304$ $1.403$ $1.218$
$\sigma_{0}$ $0.026$ $0.008$ $0.025$ $0.027$ $0.030$ $0.022$ $0.030$ $0.015$
$\sigma_{1}$ $0.030$ $0.110$ $0.204$ $0.009$ $0.259$ $0.137$ $0.364$ $0.259$
Table 3: Mean computation time (seconds) for $L_{2}$ and MLE estimation across scenarios and sample sizes.
Scenario $n=100$ $n=500$ $n=1000$ $n=5000$
$L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE $L_{2}$ MLE
BM 0.780 6.690 0.640 21.590 0.660 38.150 0.790 177.270
HT 1.510 14.320 1.690 68.750 1.630 117.780 1.540 396.950
IT 1.500 11.800 2.050 63.610 2.020 120.610 1.770 451.080
LT 1.780 12.130 2.540 66.130 2.810 133.630 3.590 687.420
Tables 1, 2, and 3 present the simulation results for median bias, MAE, and computation time, respectively.
The BM scenario exhibits the best overall performance for both estimation methods, with median bias and MAE approaching zero even at moderate sample sizes. At $n=1000$, both $L_{2}$ and MLE achieve negligible bias and MAE (all below 0.03) for all parameters. The HT scenario shows large biases at small sample sizes but rapid improvement with increasing $n$, with MLE reaching near-zero bias by $n=5000$ while $L_{2}$ converges more slowly. In contrast, the IT and LT scenarios prove substantially more challenging. For IT, both methods exhibit persistent negative bias in $\alpha$ and $\beta$ even at $n=5000$ (e.g., $L_{2}$: 0.850-0.850 and 0.469-0.469; MLE: 0.239-0.239 and 0.071-0.071). The LT scenario emerges as the most challenging, with median bias in $\beta$ exceeding 1.7 in absolute value for both methods even at the largest sample size. These difficulties arise because symmetric (IT) and highly skewed (LT) Beta distributions are harder to distinguish from the marginal distribution of $Y$, leading to weaker empirical identifiability.
Examining performance across parameters reveals distinct patterns in the relative performance of $L_{2}$ versus MLE. For the Beta shape parameters $\alpha$ and $\beta$, MLE consistently outperforms $L_{2}$ in terms of MAE, particularly in the HT and IT scenarios where MLE achieves 2 to 7 times lower MAE at $n=5000$. For the conditional mean parameters $\mu_{0}$ and $\mu_{1}$, the two methods perform more comparably in the BM scenario, while MLE shows clear advantages in the more challenging scenarios, especially for $\mu_{1}$ where the relative MAE difference can exceed an order of magnitude at large sample sizes (e.g., IT scenario at $n=5000$: 0.234 vs 0.029). The variance parameters $\sigma_{0}$ and $\sigma_{1}$ exhibit mixed results, with $L_{2}$ occasionally showing smaller bias than MLE in the BM scenario, but MLE generally achieving lower MAE in more challenging scenarios. Overall, MLE demonstrates asymptotic superiority across all parameters, though the practical advantage varies considerably by scenario and sample size, with the gap being smallest for BM and largest for IT and LT.
The computational efficiency advantage of the $L_{2}$ method is substantial and increases dramatically with sample size (Table 3). At $n=100$, $L_{2}$ is approximately 7–10 times faster than MLE across scenarios. This speedup factor increases markedly with sample size, reaching 26–41-fold at $n=500$, 48–72-fold at $n=1000$, and 192–258-fold at $n=5000$. Even in the most challenging LT scenario, $L_{2}$ estimation at $n=5000$ requires only about 3.6 seconds, on average, compared to 687 seconds for MLE—a 191-fold speedup that becomes critical when fitting models repeatedly during exploratory analysis and bootstrap inference.
@@ 6.1 Modelling apical membrane antigen 1 (AMA1) concentrations
The results of this step are presented in Figure 4. As age increases, the distribution of AMA1 antibody concentrations shifts toward higher values, indicating an overall increase in malaria exposure and infection. Another noteworthy pattern is that the bimodality, or apparent separation between the two mixture components, becomes less distinct above 15 years of age. Examining the estimates of the Beta shape parameters, $\alpha$ and $\beta$, we also observe an overall increase with age, with most of the estimate values for both parameters below 1 (indicating a U-shaped Beta) for ages below 15 years. This supports the initial interpretation based on the histograms: at younger ages, the population is more polarized, whereas with increasing age the fitted Beta distributions transition gradually from a U-shaped to a unimodal form. The increasing values of $\alpha$ and $\beta$ across age groups lead to a mean of $T$ that peaks in the $[15,20)$ interval and then gradually declines, suggesting a mild waning of antibody concentrations at older ages.
Based on the data, individuals below 15 years of age display a distinct immunological profile and antibody dynamics compared with older age groups. To illustrate how this can be accommodated, we adopt a joint modelling framework that accounts for these differences. For younger individuals, we use a mechanistic representation of $T$ based on the reversible catalytic model (11), capturing the biological processes of antibody acquisition and boosting. For older individuals, instead, we adopt a more empirical specification that flexibly describes the observed antibody distributions without imposing strong mechanistic assumptions. Furthermore, we treat the age at which this model transition occurs as an additional parameter to estimate. We also point out that this represents one of several possible analyses within our framework, and serves to demonstrate how joint modelling across ages can make fuller use of the available data, avoiding the common practice of dichotomizing or fitting separate models by age group.
@@
This constraint prevents a discontinuity in the age-dependent mean of $T$ and ensures a biologically coherent transition between the early-life dynamics of antibody acquisition and the slower antibody fluctuations observed in older individuals when the model is fitted jointly. While this continuity constraint may reduce the empirical flexibility of the fit compared to an unconstrained specification, it is imposed to retain a mechanistically interpretable parameterisation that reflects the underlying biological process.
Based on the constraints imposed on the parameters of the Beta mixture model for $T$ (see Section 3.2), an initial fit indicated that $\alpha_{1}$, $\beta_{1}$, $\beta_{2}$, and $p_{0}$ had estimates converging towards their boundary values. The fitted values were close to 1 for $\alpha_{1}$, $\beta_{2}$, and $p_{0}$, while $\beta_{1}$ resulted in a very large value of approximately $2990$. These results suggest that the first mixture component can be approximated with a Dirac measure at zero, representing individuals with very low or no seroreactivity. In other words, this finding motivates a simplified formulation in which the first component of $T$ is treated as a degenerate mass at 0, corresponding to individuals with absent or waning antibody levels, while the second component remains a continuous Beta distribution, with $\alpha_{2}$ to be estimated and $\beta_{2}=1$, representing individuals with detectable antibody responses arising from exposure to AMA1. The distribution of the latent variable $T$ for individuals below $\tau$ years of age is defined by the mixed discrete-continuous density
$f_{T}(t;a)=\begin{cases}1-\pi(a),&t=0,\\[6.0pt] \pi(a)\,\alpha_{2}\,t^{\alpha_{2}-1},&0<t<1,\\[6.0pt] 0,&\text{otherwise}\end{cases}\qquad.$ (14)
@@
Table 4 summarises the parameter estimates and uncertainty from the parametric bootstrap procedure for the joint model fitted across the full age range. The estimated change point for the transition from the mechanistic formulation to the regression-based formulation is $\hat{\tau}=20.8$ years. This suggests that the mechanistic framework remains appropriate up to approximately age 20. This finding demonstrates that rather than imposing an a priori age restriction, as is sometimes done in serology studies (e.g. yman2016), our model provides an empirical, data-driven approach to determining the appropriate age range for mechanistic modelling.
@@
The left panel of Figure 5 shows the results of the validation procedure described in Section 4.2. The main discrepancy between the fitted model and the observed data occurs in the age groups $[1,6)$ and $[6,10)$, where the model produces a distribution with a slightly higher concentration of individuals with elevated antibody levels than observed. For the remaining age groups, the observed deviations are minor. Such differences are expected because the model imposes a relatively strong structure on the age-dependent dynamics to ensure the interpretability of its components. As is often the case, a more empirical specification could yield a closer fit to the data, but at the cost of losing biological interpretability.
@@ 6.2 Modelling merozoite surface protein 1 (MSP1) concentrations
Unlike the analysis for AMA1, here we adopt a more data-driven strategy that fully exploits the flexibility of the latent-variable framework. As we will show, this enables a closer empirical fit to the MSP1 data, where we instead adopted a more mechanistically motivated approach. Based on the insights from Figure 6, we formulate an age-dependent model in which $T$ follows a Beta distribution with a smoothly varying $\alpha(a)$ and a piecewise age effect for $\beta(a)$:
@@
Table 5 summarises the parameter estimates for the MSP1 model, which point to a more durable and persistent antibody response than that observed for AMA1. The estimated exponents $\gamma$ and $\delta_{1}$ show that both $\alpha(a)$ and $\beta(a)$ increase with age, but that $\alpha(a)$ grows more rapidly, producing a steady rise in the mean latent immune state $\mathbb{E}[T]$ as given by equation (15). This rise becomes even steeper after the change point at $\zeta=11.6$ years, as reflected by the negative value of $\delta_{2}$, which slows the growth of $\beta(a)$ and therefore increases $\mathbb{E}[T]$ more sharply. The validation results in Figure 7 indicate that the model provides a good overall fit, with only minor deviations of the model-based histograms from the empirical histograms observed across the different age groups.
Figure 8 displays the expected mean of $Y$ as a function of age for both MSP1 and AMA1. AMA1 rises sharply in early childhood but then declines after about 20 years, which may reflect the higher individual-level variability in AMA1 antibody dynamics, with rapid boosting and waning following each infection (akpogheneta2008; wipasa2010; ondigo2014). The peak around 20 years may reflect the higher prevalence of asymptomatic infections with detectable parasitemia in this age range. At this stage, individuals have typically acquired clinical immunity (protection against disease symptoms) but have not yet fully developed anti-parasite immunity, sustaining infections with parasite densities high enough to continue boosting antibody levels. In contrast, as anti-parasite immunity develops in adults, lower parasite densities result in reduced antigenic stimulation and gradual antibody decline.
In contrast, MSP1 shows a steady increase across the full age range with no decline. This is consistent with MSP1 eliciting durable antibody responses (akpogheneta2008; wipasa2010). The change-point at $\zeta=11.6$ years marks an acceleration in the rate of increase. While the precise mechanism underlying this shift is unclear, plausible explanations include cumulative exposure effects associated with school attendance, immune maturation during early adolescence, or cohort effects reflecting temporal changes in malaria transmission intensity. Distinguishing among these possibilities would require longitudinal data or additional historical information on transmission patterns and intervention coverage.
@@
6.3 Model comparison and assessment of computational cost
[…]
MLE $L_{2}$
[…]
AMA1 <6<6 12,853 12,314 539 3,066.450 12,874 13,143 -269 8.128
661010 8,033 7,988 45 70.568 8,131 7,994 137 1.102
10101515 8,565 8,036 529 37.449 8,565 8,040 525 0.842
15152020 4,946 4,470 476 12.213 4,946 4,472 474 1.366
20203030 6,016 5,316 700 13.675 6,017 5,318 699 0.666
30304545 5,084 4,300 784 9.989 5,085 4,301 784 1.108
$45+$ 6,525 5,131 1,394 11.018 6,526 5,133 1,393 0.857
MSP1 <6<6 13,277 13,235 42 1,181.518 13,328 13,291 37 7.540
661010 7,816 7,654 162 61.514 7,816 7,674 142 0.750
10101515 8,678 8,364 314 33.332 8,679 8,382 297 1.617
15152020 5,184 4,973 211 17.352 5,185 4,980 205 1.255
20203030 6,129 5,870 259 19.832 6,129 5,878 251 1.238
30304545 5,071 4,843 228 13.780 5,071 4,849 222 1.169
$45+$ 6,391 5,932 459 14.386 6,391 5,941 450 1.244
We conduct an age-stratified analysis to compare the proposed latent Beta model (LBM) with the standard Gaussian mixture model (GMM). For the LBM, we assume a single Beta distribution for the latent variable $T$, with shape parameters $\alpha$ and $\beta$ to be estimated. For both models, the conditional distribution parameters $\mu_{0}$, $\mu_{1}$, $\sigma_{0}$, and $\sigma_{1}$ are fixed at values estimated by fitting the GMM and LBM, respectively, to the youngest age group (<6<6 years), which shows the clearest separation between individuals with low and high antibody levels. For each subsequent age group, we re-estimate only the mixing proportion $\pi$ in the GMM or the Beta shape parameters $\alpha$ and $\beta$ in the LBM. Since the GMM can be viewed as a special case of the LBM in which the latent distribution collapses to a discrete two-point distribution, this constrained comparison directly assesses whether age-related variation in antibody responses can be explained through changes in the latent distribution structure for the LBM, as opposed to changes in mixture weights for the GMM.
The results of this comparison are presented in Table 6. For LBM the fitted densities using MLE are shown in the left panel of Figure 4 and Figure 6; for GMM these are shown in Figure 9 and Figure 10 of the Appendix. Under maximum likelihood estimation, the LBM consistently outperforms the GMM across all age groups and both antigens. For AMA1, the BIC advantage of the LBM increases with age and becomes substantial from age 10 years onwards, often exceeding 100 BIC units in difference. For MSP1, the BIC differences in favour of the LBM are generally smaller than those observed for AMA1, indicating that the GMM is better able to adjust to age-related changes in MSP1 antibody distributions than for AMA1. Nevertheless, the LBM still provides a superior fit across all age groups, suggesting that a continuous latent representation remains more appropriate even when the Gaussian mixture adapts relatively well.
When using the $L_{2}$ histogram-based optimisation, the qualitative pattern of results remains largely unchanged. The only exception is the youngest age group for AMA1, where the $L_{2}$ criterion slightly favours the GMM, reflecting the fact that both models rely on parameters fixed using this age group and that the $L_{2}$ approximation places greater emphasis on matching the empirical histogram. Aside from this case, the BIC differences obtained under $L_{2}$ optimisation are highly consistent with those derived from maximum likelihood, both in direction and magnitude.
From a computational perspective, the $L_{2}$ histogram-based optimisation yields substantial efficiency gains relative to maximum likelihood estimation. For the LBM, speedups range from more than an order of magnitude for moderate-sized age groups to several hundred-fold for the largest groups, reducing computation times from tens of minutes to a few seconds. These gains are particularly pronounced for AMA1, but are also substantial for MSP1. Importantly, this dramatic reduction in computational cost is achieved without altering our conclusions, indicating that the $L_{2}$ approach provides a reliable and efficient approximation for large-scale model comparisons.